# Who Gets Cited Most? Benchmarking Long-Context Reasoning on Scientific Articles

### v2

@@ 6 Conclusion
We introduce SciTrek, a benchmark for evaluating LLMs’ ability to perform multi-document information synthesis and structured reasoning over full-text scientific articles. By generating questions and ground-truth answers through a SQL backbone operating on article metadata, we provide a highly scalable framework with explicit reasoning processes that enables fine-grained error analysis.
Our extensive evaluation demonstrates that SciTrek poses a significant challenge to open-weight and proprietary LLMs, with only modest performance gains observed through supervised fine-tuning and reinforcement learning. Specifically, models struggle with compound logical conditions, fail at sorting tasks, and often produce incomplete or poorly formatted outputs. We believe the proposed methodology for creating SciTrek generalizes beyond titles, authors, and references to encompass broader metadata elements and domains, offering a tool to diagnose persistent shortcomings in LCLMs and improve their capability to support scientific workflows.
@@ 7 Limitations
While SciTrek benchmarks fundamental reasoning capabilities (including filtering, sorting, aggregating and their combinations) of LCLMs on scientific articles, it covers only a limited set of core elements of scientific articles, specifically titles, authors, and references. Future work will explore additional long-context model capabilities for understanding scientific articles, such as interpreting figures and performing domain-specific content reasoning.
Appendix A Details on SciTrek Construction
@@
Database Column Data
Table Name Type Description
@@
Table 10: Description of the database tables that we use to curate SQL queries and answers.
@@
Table 11: Distribution of SQL commands and operators in our test data. $\%$ refers to the prevalence of each command/operator in our test data’s SQL queries. These percentages do not need to add to 100%.
Appendix B Experimental Settings
Configuration details for the models we use are presented in Table 12. For all models, we use their default inference configuration for answer generation from their Huggingface repositories. We include both reasoning models and non-reasoning models. Qwen3-4B-Thinking-2507, Qwen3-30B-A3B-Thinking-2507, DeepSeek-R1-Distill-Llama-70B, Gemini 2.5 Pro and o4-mini are reasoning models. In our experiments, o4-mini conducts reasoning with the default thinking effort medium, Gemini 2.5 Pro was given a thinking budget of 512 tokens, and Qwen3-4B-Thinking-2507, Qwen3-30B-A3B-Thinking-2507, and DeepSeek-R1-Distill-Llama-70B were prompted with the reasoning prompt in Figure 6.
@@ Appendix B Experimental Settings
Table 12: Proprietary and open-weight LCLMs that are evaluated in our experiments. (Parameter numbers for proprietary models are not accessible. “17B×16E" indicates that the model uses a mixture-of-experts architecture with 16 experts, each containing 17 billion parameters. Context length is based on the maximum number of tokens these models can handle as input.)
@@
Table 13: Evaluation of open-weight (top) and proprietary (bottom) models on the SciTrek benchmark using F1 (%). Results are reported for two settings: using the full-text scientific articles as context and using the corresponding database tables. F-* refers to test data showing full-text article clusters with different maximum lengths, 64K/128K/512K/1M. D-* refers to test data showing the database tables corresponding to F-64/128/512/1024 as input. Note that for any given article cluster, the database representation is significantly shorter (mean length $\sim$}2K tokens); therefore we are able to pass database representations for all lengths to all models. Context Size denotes the maximum context length supported by each model. — indicates the model cannot handle the given context size; $\star$ indicates that the model was not evaluated due to prohibitive computational cost.
Appendix C Additional Results Using F1
@@
Table 13 presents additional results using F1 as the evaluation metric. Models are tested against full-text scientific articles and database tables as context. The input length averages only 1,980 tokens when database tables of are used as context across different lengths of full-text article collections.
Appendix D Details on Model Failure Analysis
To gain deeper insight into model behaviour, we analyzed the failure patterns exhibited in model-generated answers on SciTrek. We randomly selected 60 instances in the input length of 128K that were hard for all zero-shot and supervised models (10 instances per skill in Table 4). We find that models tend to output “NULL" particularly on samples that require the skill of filtering. As shown in Table 14, most open-weight models tend to randomly output “NULL" for any sample. This indicates that these models rely more on the instruction to produce a fallback “NULL” answer rather than on genuine understanding of the provided context. [^15] Our instruction prompted models to generate “NULL” when they could not get the answer from the context. In addition, we find that models fail to follow the specified format for the answers, especially for sorting-related questions, as shown in Table 15. For example, models tend to generate author lists when sorted author numbers are requested and output lists when aggregates are requested. Interestingly, GPT-4.1 produces incorrectly formatted answers far more frequently than all other models. We further observe that some models tend to produce partial answers when lists are requested, especially for tasks that require aggregation. As shown in Table 16, our GRPO-based model tends to generate partial answers more than other models.
We also analyzed model performance on questions involving negation (e.g., those containing adverbs such as “not” or “never”). For questions that involve filtering-related skills, we have 132 instances with negation in our test set (Filtering: 11, Filtering+Aggregation: 15, Filtering+Sorting:12, and Relational Filtering:94). The results in Table 17 show that all models struggle with negation-based filtering.
@@ Appendix D Details on Model Failure Analysis
We further manually analyzed 200 reasoning traces from reasoning models including Qwen2.5-7B-1M (GRPO), Qwen3-4B-A3B-Thinking-2507 and DeepSeek-R1-Distill-Llama-70B for 25 random SQL query templates. We cannot obtain reasoning traces from Gemini 2.5 Pro, GPT-4.1 and o4-mini for analysis. For each SQL query template, we have different SQL queries and questions for different sizes of contexts, while the questions require the same reasoning capability. There are no reasoning traces when the input is 512K or 1M tokens for Qwen3-4B-Thinking-2507 and DeepSeek-R1-Distill-Llama-70B, as they do not support the input of 512K or 1M tokens.
Model A S F F+A F+S RF
Qwen2.5-7B-Instruct-1M 70 90 80 90 80 80
Qwen2.5-14B-Instruct-1M 10 10 20 50 50 60
Gemma-3-27B-IT 10 0 0 40 10 40
Llama-4-Scout-14Bx16E-Instruct 30 20 30 60 20 60
Llama-3.3-70B-Instruct 10 10 30 80 50 50
DeepSeek-R1-Distill-Llama-70B 0 0 0 10 0 10
Gemini 2.5 Pro 0 0 0 20 30 10
GPT-4.1 0 0 0 10 0 0
o4-mini 0 0 10 10 30 20
Qwen2.5-7B-Instruct-1M (SFT) 0 0 40 30 40 0
Qwen2.5-7B-Instruct-1M (GRPO) 0 0 0 0 0 0
Table 14: Proportion of samples (%) where models output “NULL” broken down across information processing skills. Zero-shot models are shown in the first block, supervised versions of Qwen2.5-7B-Instruct-1M are shown in the second block. A abbreviates Aggregation, S stands for Sorting, F for Filtering, and RF for Relational Filtering.
Model A S F F+A F+S RF
Qwen2.5-7B-Instruct-1M 0 0 0 0 0 0
Qwen2.5-14B-Instruct-1M 0 30 10 0 30 0
Gemma-3-27B-IT 0 20 30 20 20 0
Llama-4-Scout-14Bx16E-Instruct 10 10 0 10 30 10
Llama-3.3-70B-Instruct 10 10 30 0 10 0
DeepSeek-R1-Distill-Llama-70B 0 0 0 0 0 0
Gemini 2.5 Pro 0 10 0 0 0 0
GPT-4.1 30 30 20 50 20 0
o4-mini 10 20 0 0 0 0
Qwen2.5-7B-Instruct-1M (SFT) 0 0 0 0 0 0
Qwen2.5-7B-Instruct-1M (GRPO) 0 0 0 10 0 0
Table 15: Proportion of samples (%) where models do not follow the specified answer format broken down across information processing skills. Zero-shot models are shown in the first block, supervised versions of Qwen2.5-7B-Instruct-1M are shown in the second block. A abbreviates Aggregation, S stands for Sorting, F for Filtering, and RF for Relational Filtering.
Model A S F F+A F+S RF
Qwen2.5-7B-Instruct-1M 10 0 10 0 0 0
Qwen2.5-14B-Instruct-1M 0 0 0 0 0 10
Gemma-3-27B-IT 10 10 10 0 0 0
Llama-4-Scout-14Bx16E-Instruct 0 0 0 0 0 0
Llama-3.3-70B-Instruct 0 0 10 0 0 0
DeepSeek-R1-Distill-Llama-70B 0 0 0 20 0 0
Gemini 2.5 Pro 20 0 10 20 0 10
GPT-4.1 10 0 0 0 0 0
o4-mini 10 0 20 30 10 0
Qwen2.5-7B-Instruct-1M (SFT) 10 0 0 0 0 20
Qwen2.5-7B-Instruct-1M (GRPO) 20 0 30 10 20 0
Table 16: Proportion of samples (%) where models generate partial answers broken down across information processing skills. Zero-shot models are shown in the first block, supervised versions of Qwen2.5-7B-Instruct-1M are shown in the second block. A abbreviates Aggregation, S stands for Sorting, F for Filtering, and RF for Relational Filtering.
Model F+A F+S F RF Avg
Qwen2.5-7B-Instruct-1M 0 0 0 0 0
Qwen2.5-14B-Instruct-1M 0 20.0 0 2.1 3.8
Gemma-3-27B-IT 0 6.7 0 0 0.8
Llama-4-Scout-14Bx16E-Instruct 0 20.0 0 2.1 3.8
Llama-3.3-70B-Instruct 0 6.7 0 0 0.8
DeepSeek-R1-Distill-Llama-70B 0 13.3 0 0 1.5
Gemini 2.5 Pro 0 20.0 8.3 6.4 9.1
GPT-4.1 0 13.3 0 2.1 3.8
o4-mini 0 73.3 41.7 7.4 22.0
Qwen2.5-7B-Instruct-1M (SFT) 0 6.7 0 11.7 9.1
Qwen2.5-7B-Instruct-1M (GRPO) 0 33.3 8.3 3.2 8.3
Table 17: Model performance on questions that involve negation (e.g., “not" or “never"), using exact match (%). Zero-shot models are shown in the first block, supervised versions of Qwen2.5-7B-Instruct-1M are shown in the second block. A abbreviates Aggregation, S stands for Sorting, F for Filtering, and RF for Relational Filtering.
Appendix E Prompts used for SciTrek Creation and Model Evaluation
@@
We generate natural language questions by prompting Qwen2.5-Coder-32B to convert SQL queries to natural language. Our prompt is presented in Figure 3. The prompt that we use to convert natural language questions back to to SQL queries is shown in Figure 4. We employ two different prompts for (zero-shot) model evaluation. The prompt for assessing LCLM capabilities against full-text articles is presented in Figure 5, and the prompt for using the corresponding database tables as context is shown in Figure 7. The prompt used by GRPO to generate reasoning races is given in Figure 6.

### v3

@@ 6 Conclusion
We introduce SciTrek, a scalable benchmark for diagnosing long-context numerical reasoning in LLMs. Questions and ground-truth answers are generated automatically via SQL queries over article metadata, providing verifiable reasoning traces that enable fine-grained error analysis without manual annotation.
Our experiments reveal a pronounced gap between performance on structured database tables and full-text articles, a gap that directly quantifies the cost of reasoning over unstructured long contexts and exposes systematic failure modes: miscounting bibliographic elements, misinterpreting compound logical conditions (particularly negation), and producing incomplete or incorrectly formatted outputs. Despite modest improvements from supervised fine-tuning and reinforcement learning, post-training on SciTrek does transfer to broader long-context benchmarks (LongBench v2 and Loong), suggesting that the numerical-reasoning skills it targets are genuinely general. Analysis of reasoning traces further shows that GRPO strengthens abstract multi-step reasoning while leaving fine-grained numerical operations, most notably reference counting, largely unresolved.
We believe the proposed methodology generalises beyond titles, authors, and references to richer metadata (e.g., patient records, codebases, news archives) and other domains in which entities can be structured as relational tables, making it straightforward to extend as model context windows continue to expand.
@@ 7 Limitations
SciTrek provides a useful diagnostic benchmark for evaluating the numerical reasoning capabilities of long-context language models, but several limitations and areas for improvement are worth noting:
Automatically generated questions.
Our questions are generated by an LLM and target surface-level article metadata rather than domain-specific content. As a result, they do not capture the full complexity of real-world queries, which may require pragmatic inference, implicit comparisons, or background domain knowledge. That said, the benchmark deliberately targets fundamental operations — counting, filtering, aggregating, sorting — that are prerequisite capabilities for more complex analytical tasks.
Coverage of scientific domains.
All article collections are drawn from arXiv-adjacent fields. Whether the observed failure patterns generalise to other scientific disciplines, or to non-scientific long-context settings, remains an open question. Future work could extend the benchmark to additional disciplines and modalities (e.g., tables and figures).
Evaluation metrics.
Exact match is strict: a sorted list with a single incorrect element receives a score of zero, regardless of how close the answer is. F1 partially mitigates this, but neither metric handles semantically equivalent reformulations, such as abbreviated author names or minor title variations. More nuanced evaluation metrics for structured outputs remain an open problem.
Coverage of metadata elements.
SciTrek currently covers only titles, authors, and references. Richer elements such as experimental results, datasets, and cited methods are not yet included. Extending the benchmark to encompass these would enable evaluation of deeper analytical tasks over scientific articles.
Appendix A Numerical Reasoning in Expert-Written Benchmark Questions
Our numerical reasoning questions are constructed around atomic SQL operations, specifically sorting, filtering, and aggregating, and operators covering comparsion, arithmetic and logical reasoning, shown in Table 3. These capabilities are fundamental building blocks of complex, realistic tasks. QMDR (Wolfson et al., 2020) provides a widely used formalism, in which complex questions can be decomposed into atomic sub-questions that require filtering, aggregating, comparing, sorting, logical and arithmetic operations. More recently, MoNaCo (Wolfson et al., 2025) demonstrates through manual decomposition of naturally occurring human-written questions that real-world complex questions in multi-document understanding frequently demand aggregation and arithmetic skills.
Our manual analysis of human-written or human-audited benchmark questions in OpenScholar (Asai et al., 2024), LongBench v2 (Bai et al., 2025), and Loong (Wang et al., 2024c) reveals that these questions implicitly require the same numerical reasoning capabilities targeted by SciTrek. Example questions are shown in Table 15. Notably, such capabilities arise even in expert-written financial questions from LongBench v2, underscoring their relevance well beyond scientific text.
Index Question Reasoning
Q1-OS "What are the latest works on finetuning an auto-regressive LM for dense passage retrieval? How are their performance compared with bi-directional encoders?" sorting, arithmetic, logical
Q2-OS "Which downstream task can solved by AlphaFold3 but cannot performed by ESM-3?" filtering, logical
Q3-OS “Compared to 2023, how has the percentage of finished goods apparel factories from countries other than Vietnam, China, and Cambodia changed in 2024?” arithmetic, aggregating, comparing
Q4-LB “In the financial reports of Apple Inc. and Samsung Electronics for the years 2022 and 2023, which company has a higher percentage of revenue derived from the product category of phones, and in what range do the differences in this dependency between the two companies in the two years fall?” comparing, filtering, aggregating, logical, arithmetic
Q5-LG We kindly ask you to thoroughly review the provided papers and construct a citation chain from them. …, and ensure that the citation chain accurately reflects the sequential citation order among these documents. …" sorting, filtering
Table 15: Example questions from OpenScholar (OS), LongBench v2 (LB), and Loong (LG). These questions implicitly require numerical reasoning targeted by SciTrek.
Appendix B Details on SciTrek Construction
@@
Database Table Column Name Data Type Description
@@
Table 16: Description of database tables used to curate SQL queries and answers.
@@
Table 17: Distribution of SQL commands and operators in our test data. Percentages reflect prevalence of each command or operator among SQL queries and need not add to 100%.
Appendix C Experimental Settings
Configuration details for all models are presented in Table 18, with default inference settings from their Huggingface repositories used throughout. Our evaluation covers both reasoning and non-reasoning models. The reasoning models are Qwen3-4B-Thinking-2507, Qwen3-30B-A3B-Thinking-2507, DeepSeek-R1-Distill-Llama-70B, Gemini 2.5 Pro, and o4-mini, which differ in their thinking configurations: o4-mini uses the default medium thinking effort, Gemini 2.5 Pro is allocated a budget of 512 tokens, and the three open-source models were prompted with the reasoning prompt in Figure 6.
@@ Appendix C Experimental Settings
Table 18: Proprietary and open-weight LLMs evaluated in our experiments. (Parameter counts are unavailable for proprietary models. “17B×16E" denotes a mixture-of-experts architecture with 16 experts, each containing 17 billion parameters. Context length refers to the maximum number of input tokens supported by each model.)
@@
Table 19: Evaluation SQL execution-based method (top), open-weight (middle) and proprietary (bottom) models on SciTrek using F1 (%). Results are reported for two context settings: full-text scientific articles and corresponding database tables. F-* denotes test instances using full-text article collections up to 64K, 128K, 512K, and 1M tokens. D-* denotes test instances using the database tables corresponding to F-64, F-128, F-512, and F-1024. For any given article cluster, the database representation is significantly shorter (mean length $\sim$}2K tokens) and can therefore be passed to all models regardless of their context limit. Context Size denotes the maximum input length supported by each model. — indicates the model cannot process the given context size; $\star$ indicates the model was not evaluated due to prohibitive computational cost.
Appendix D Prompt Templates for Extracting Titles, Authors, and References
We use LLMs to extract titles, authors, references from article texts; the corresponding prompts are shown in Table 20. For references specifically, we use two separate prompts: one to count references and one to detect the citation relationships between any two scientific articles, rather than extracting and outputting the full reference list directly. Both operations are applied to article content in chunks of 16K tokens, as the extraction model (Qwen2.5-Coder-32B-Instruct in our experiments) does not reliably handle longer contexts.
The Instruction for Title Extraction
You are given a chunk of an scientific article.
<CONTEXT>
If the article title is included in the given text, please locate and extract the title.
You must output the title in the following JSON format, and do not include any extra text.
{"title": "the extracted title"}
If there is no article title in the given text, just output {"title": ""}
The Instruction for Authors Extraction
You are given a chunk of an scientific article.
<CONTEXT>
If the author list of the article is included in the given text, please locate and extract the author list.
You must output the author list in the following JSON format, and do not include any extra text.
{"authors": "the author list separated with commas"}
If there is no author list in the given text, just output {"authors": ""}
The Instruction for Counting References
You are given a chunk of an scientific article.
<CONTEXT>
If an integral or segmental reference section is included in the given text, please count the number of the references.
You must output the reference count in the following JSON format, and do not include any extra text.
{"reference_count": "the number of references"}
If there is no any reference section in the given text, just output {"reference_count": "0"}
The Instruction to Detect Citation Relationships
You are given the title of one scientific article (Article A) and a chunk of another scientific article (Article B).
The title of Article A:
<TITLE>
The chunk of Article B:
<CONTEXT>
Please check whether Article A is cited by Article B. In other words, the task is to check if the title of Article A exists in the given chunk of Article B.
You must output your answer in the following JSON format, and do not include any extra text.
{"a_cited_by_b": "true or false"}
Table 20: Prompt templates for extracting titles, authors, and references from article text. <CONTEXT> and <TITLE> are placeholders for the input article chunk and the article title, respectively.
Appendix E Additional Results Using F1
@@
Table 19 reports additional results using F1 as the evaluation metric. Models are evaluated with either full-text scientific articles or database tables as context. When database tables are used, the the average input length is only 1,980 tokens, regardless of the size of the underlying article collection.
Appendix F Additional Details on Reasoning Trace Analysis
@@
We manually analyzed 200 reasoning traces from reasoning models, including Qwen2.5-7B-1M (GRPO), Qwen3-4B-A3B-Thinking-2507 and DeepSeek-R1-Distill-Llama-70B for 25 random SQL query templates. We cannot obtain reasoning traces from Gemini 2.5 Pro, GPT-4.1 and o4-mini for analysis. For each SQL query template, we have different SQL queries and questions for different sizes of contexts, while the questions require the same reasoning capability. There are no reasoning traces when the input is 512K or 1M tokens for Qwen3-4B-Thinking-2507 and DeepSeek-R1-Distill-Llama-70B, as they do not support the input of 512K or 1M tokens.
Appendix G Prompts used for SciTrek Creation and Model Evaluation
@@
We generate natural language questions by prompting Qwen2.5-Coder-32B to convert SQL queries into natural language; the corresponding prompt is shown in Figure 3. The prompt for the reverse conversion, from natural language questions back to SQL queries, is shown in Figure 4. For zero-shot model evaluation, we use two separate prompts: one for assessing LLM capabilities against full-text articles (Figure 5) and one for using the corresponding database tables as context (Figure 7). Finally, the prompt used by GRPO to generate reasoning traces is given in Figure 6.